# Hidden yet quantifiable: A lower bound for confounding strength using randomized trials

### v3

@@ 1 Introduction
Monitoring the performance of a newly approved treatment is crucial, a process commonly referred to as post-marketing surveillance [62]. Nowadays, the U.S. Food and Drug Administration promotes the integration of observational data in this process to address the shortcomings of randomized evidence [48, 39]. This strategy is essential for validating personalized treatments, like immunotherapy for certain types of cancer, where randomized evidence is scarce, and treatment costs are substantial [25, 23].
Yet, unobserved confounding can significantly compromise causal conclusions drawn from observational data. To tackle this issue, sensitivity analysis has been the prevalent paradigm since its conception by Cornfield et al. [10]. This field studies how a specific strength of unobserved confounding affects causal conclusions and introduces the concept of a critical value [59, 33], i.e. the minimum strength unobserved confounders would need to have to explain away the estimated treatment effect. However, critical values are solely based on observational data and can differ substantially from the true confounding strength. As a result, epidemiologists often rely on heuristic judgments to decide whether an observational study is flawed.
Estimating the true confounding strength is infeasible without further assumptions. Yet, once a treatment gains approval, we may have access to a randomized trial that allows for more effective strategies to address unobserved confounding. A recent line of works proposes to combine the estimators from randomized and observational data, e.g. see Colnet et al. [7], Brantner et al. [3] for a survey. However, these methods crucially rely on some prior knowledge of the confounding bias structure, that is not always available in practice.
@@
We propose an alternative strategy to leverage randomized trials, that is, to test and quantify the true confounding strength. In particular, if strong confounding is detected, epidemiologists can take proactive measures to correct it. Most directly, they can identify and incorporate relevant covariates into the study design if they were initially overlooked [19]. On the other hand, if small confounding is detected, epidemiologists can continue their analysis (see Figure 1 for an illustration of the pipeline). More concretely, our contributions are as follows.
@@ 1.1 Related work
Our approach is closely related to a line of work that proposes statistical tests for the presence of unobserved confounding. In particular, several works leverage randomized trials to detect unobserved confounding. These tests check for significant differences between average treatment effect estimates obtained from randomized and observational data [61, 65, 42, 30]. More sophisticated approaches also test for differences in conditional average treatment effect estimates [31] and account for right-censored outcomes [15].
Similarly, other works have designed statistical tests using instrumental variables and negative control outcomes instead of randomized trials [40, 12, 16, 53]. Additionally, multiple observational studies can be leveraged to test conditional independences and detect unobserved confounding [36].
@@
Finally, another line of works proposes calibrating the value of confounding strength using only observational data [29, 60]. However, the true confounding strength can be arbitrarily different from the calibrated strength, and there is no theoretical result for how these two quantities are related, even with infinite samples.
@@ 2 Setting and notation
where $\mathbb{P}_{\operatorname{det}}$ is deterministically given by $Y=Y(T)$ [^3] Given that samples are drawn i.i.d., this asumption is equivalent to the classic SUTVA [51]., $\mathbb{P}_{\operatorname{inv}}$ is invariant across studies, and $\mathbb{P}_{\operatorname{cnf}}^{\diamond}$ differs for $\diamond\in\{\mathrm{rct},\mathrm{os}\}$. This factorization captures the essence of the potential outcome framework, where $Y(1)$ and $Y(0)$ do not depend on $T$ while being more general. In particular, it allows for shifts in the marginal distribution of the observed and unobserved confounders.
We illustrate the corresponding graphical model in Figure 2. Note that numerous attempts have been made to unify potential outcomes and graphical models, with the most prominent being the Single World Intervention Graphs [50]. However, we propose a simpler graphical model in this context since we do not use the graph to infer counterfactual independencies.
@@
This property is standard for generalizing the findings of randomized trials to another population [46, 7, 13], and is a weaker assumption than ignorability of study selection [56, 28] or sample ignorability of treatment effects [38].
@@
This assumption is strictly weaker than the positivity of trial participation [54, 24, 2, 43, 8]. It is also expected to hold in our setting, as it aligns with the design of observational studies by regulatory agencies for drug monitoring  [26], and particularly for post-marketing surveillance [20, 52].
@@ 2.1 Sensitivity analysis
Sensitivity analysis is commonly used to account for unobserved confounding in observational data. In particular, this approach estimates an interval for the treatment effect that depends on an assumed confounding strength $\Gamma$ of $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$. Throughout the paper, we define the confounding strength using the widely accepted marginal sensitivity model [57].
@@
Under this notion of confounding strength, we can define a set of full distributions ~$\tilde{\mathbb{P}}_{\operatorname{full}}$ that are compatible with the marginal distribution of the observational study $\mathbb{P}^{\operatorname{os}}$ and have a bounded odds ratio.
@@
Given a distribution $\mathbb{P}^{\operatorname{os}}$ over $(X,Y,T)$ and a confounding strength $\Gamma\geq 1$, we define the set $\mathcal{E}(\mathbb{P}^{\operatorname{os}},\Gamma)$ of distributions ~$\tilde{\mathbb{P}}_{\operatorname{full}}$, as
@@
Above, we do a slight abuse of notation by defining $\mu$ as both a function and a real number, depending on its argument. Several estimators have recently emerged in the literature for the CATE bounds [35, 32, 47] and for the ATE bounds [66, 18, 17]. We will leverage these estimators to construct a statistical test that detects unobserved confounding above a certain strength.
@@ 3.1 Statistical tests for $H_{0}(\Gamma)$
In what follows, we have access to a randomized trial $D_{\operatorname{rct}}=\{(X_{i},Y_{i},T_{i})\}_{i=1}^{n_{\operatorname{rct}}}$ sampled i.i.d from the distribution $\mathbb{P}^{\operatorname{rct}}$, and an observational study $D_{\mathrm{os}}=\{(X_{i},Y_{i},T_{i})\}_{i=1}^{n_{\operatorname{os}}}$, sampled i.i.d. from the distribution $\mathbb{P}^{\operatorname{os}}$. We first propose estimates for the average treatment effect under two target populations. Then, we leverage these estimates together with the sensitivity bounds to design an asymptotically valid statistical test at significance level $\alpha$. Finally, we show how such a test can be used to establish an asymptotically valid lower bound on the unobserved confounding strength.
@@
$\mu(\mathbb{P}^{\diamond}_{X},\mathbb{P}^{\diamond}_{\operatorname{full}})$ $=\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[Y\left(\frac{T}{\pi}-\frac{(1-T)}{1-\pi}\right)w(X)\right],\quad\mathrm{where}\quad w(X):=\frac{\mathbb{P}^{\diamond}(X)}{\mathbb{P}^{\operatorname{rct}}(X)}.$
Lemma 3.2 is a well-known result in the transportability literature [6, 9]. Essentially, it establishes that when the distribution shift between $\mathbb{P}^{\operatorname{rct}}_{X}$ and $\mathbb{P}^{\diamond}_{X}$ can be corrected, we can identify and estimate the ATE under $\mathbb{P}^{\diamond}_{X}$.
@@
  • For $\mathbb{P}^{\diamond}_{X}=\mathbb{P}^{\operatorname{rct}}_{X}$, we estimate the CATE sensitivity bounds from observational data and average them over the target population. Specifically, we use the B-Learner [47] to estimate the sensitivity analysis bounds.
  • For $\mathbb{P}^{\diamond}_{X}=\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}$, we have two options: either estimate the CATE sensitivity analysis bounds and average them, or directly estimate the ATE sensitivity analysis bounds over the target population. In our experiments, we directly estimate the ATE sensitivity analysis bounds using either the DVDS [18] or the QB estimator [17].
@@
Let $\hat{\phi}_{\diamond}(\Gamma,\alpha)$ be the test defined in Algorithm 1, for a fixed $\Gamma\in[1,\infty)$ and significance level $\alpha$. Then, under Assumptions 2.12.3 and the setting described in Section 2, we have, for $H_{0}(\Gamma)$,
@@
$\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)$ is a valid asymptotic test at level $\alpha$.
(ii) If $\hat{\mu}_{\Gamma}^{+}$ and $\hat{\mu}_{\Gamma}^{-}$ are consistent estimators of the ATE sensitivity analysis bounds that satisfy
$\sqrt{n_{\operatorname{os}}}\;\hat{\mu}_{\Gamma}^{+}\xrightarrow{\mathcal{D}}\mathcal{N}(\mu_{\Gamma}^{+},(\sigma_{\Gamma}^{+})^{2}),\;\sqrt{n_{\operatorname{os}}}\;\hat{\mu}_{\Gamma}^{-}\xrightarrow{\mathcal{D}}\mathcal{N}(\mu_{\Gamma}^{-},(\sigma_{\Gamma}^{-})^{2}),$
$\hat{\phi}_{\widetilde{\mathrm{os}}}(\Gamma,\alpha)$ is a valid asymptotic test at level $\alpha$.
We provide a complete proof in Section A.1.2. Notably, Assumption (ii) is relatively mild and expected to hold for various estimators; for instance, it can be satisfied by the DVDS estimator [18]. On the other hand, Assumption (i) is stronger and generally only expected to hold when $n_{\operatorname{os}}\gg n_{\operatorname{rct}}$.
In essence, we propose two tests that work under different assumptions: $\hat{\phi}_{\mathrm{rct}}$ relies on a consistent estimate of the CATE sensitivity analysis bounds, while $\hat{\phi}_{\widetilde{\mathrm{os}}}$ requires an estimate of the importance weights $w(X)$ [^4] The importance weights can be identified when the observational study and the randomized trial adhere to a nested trial design [44, 45, 5]. See Section A.2 for a discussion on how the importance weights are estimated in this setting..
@@
$\hat{\mu}_{\Gamma}^{+}=\hat{\mathbb{E}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{+}_{\Gamma}(X)],\quad(\hat{\sigma}^{+}_{\Gamma})^{2}=\widehat{\operatorname{Var}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{+}_{\Gamma}(X)],\quad\hat{\mu}_{\Gamma}^{-}=\hat{\mathbb{E}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{-}_{\Gamma}(X)],\quad(\hat{\sigma}^{-}_{\Gamma})^{2}=\widehat{\operatorname{Var}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{-}_{\Gamma}(X)],$
where $\hat{\mathbb{E}}[\cdot]$ and $\widehat{\operatorname{Var}}[\cdot]$ denote the empirical mean and variance, respectively.
@@ 3.2 A lower bound on unobserved confounding strength
Let $\hat{\Gamma}_{\texttt{LB}}$ be as in Equation (2) for a fixed significance level $\alpha$. Then, under Assumptions 2.12.3 and the setting described in Section 2, $\hat{\Gamma}_{\texttt{LB}}$ is an asymptotically valid lower bound, i.e.
$\mathbb{P}(\hat{\Gamma}_{\texttt{LB}}\leq\Gamma^{\star})\geq 1-\alpha-o_{\mathbb{P}}(1).$
@@
$\leq\mathbb{P}(\hat{\phi}_{\mathrm{rct}}(\Gamma^{\star},\alpha)=1)\leq\alpha+o_{\mathbb{P}}(1),$
@@ 4 Synthetic Experiments
Intuitively, the sensitivity analysis bounds are tight for a specific $\Gamma^{\star}$ when $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$ leads to a marginal distribution $\mathbb{P}^{\operatorname{os}}$ that maximally biases the estimable ATE. This situation occurs, for instance, when patients experiencing smaller outcomes are assigned to the control group while those with larger outcomes are in the treatment group. In this case, the sensitivity analysis bounds must be sufficiently large to include the true ATE and remain valid. Such a scenario is only possible if $U$ is very informative of $Y(1)$, captured by a high correlation coefficient. Conversely, when $Y(1)$ is independent of $U$, the true ATE is unaffected by the unobserved confounding, and the sensitivity bounds are unnecessarily conservative, leading to low power of the test and hence looser $\hat{\Gamma}_{\texttt{LB}}$. More formally, in Section A.3 we show that when the confounder $U$ is equal to $(Y(1),Y(0))$, the correlation coefficient $\rho_{u,y}=1$ and $\hat{\Gamma}_{\texttt{LB}}$ converges to $\Gamma^{\star}$ in the infinite sample limit. In contrast, when $U$ is independent of $(Y(1),Y(0))$, $\rho_{u,y}=0$ and $\hat{\Gamma}_{\texttt{LB}}=1$.

### v4

@@ 1 Introduction
Monitoring the performance of a newly approved treatment is crucial, a process commonly referred to as post-marketing surveillance [vlahovic2011postmarketing]. Nowadays, the U.S. Food and Drug Administration promotes the integration of observational data in this process to address the shortcomings of randomized evidence [platt2018fda, klonoff2020new]. This strategy is essential for validating personalized treatments, like immunotherapy for certain types of cancer, where randomized evidence is scarce, and treatment costs are substantial [hayden2016gene, goetz2018personalized].
Yet, unobserved confounding can significantly compromise causal conclusions drawn from observational data. To tackle this issue, sensitivity analysis has been the prevalent paradigm since its conception by cornfield59. This field studies how a specific strength of unobserved confounding affects causal conclusions and introduces the concept of a critical value [vanderweele2017sensitivity, jin2023sensitivity], i.e. the minimum strength unobserved confounders would need to have to explain away the estimated treatment effect. However, critical values are solely based on observational data and can differ substantially from the true confounding strength. As a result, epidemiologists often rely on heuristic judgments to decide whether an observational study is flawed.
Estimating the true confounding strength is infeasible without further assumptions. Yet, once a treatment gains approval, we may have access to a randomized trial that allows for more effective strategies to address unobserved confounding. A recent line of works proposes to combine the estimators from randomized and observational data, e.g. see colnet2020causal, brantner2023methods for a survey. However, these methods crucially rely on some prior knowledge of the confounding bias structure, that is not always available in practice.
@@
We propose an alternative strategy to leverage randomized trials, that is, to test and quantify the true confounding strength. In particular, if strong confounding is detected, epidemiologists can take proactive measures to correct it. Most directly, they can identify and incorporate relevant covariates into the study design if they were initially overlooked [dreyer2018advancing]. On the other hand, if small confounding is detected, epidemiologists can continue their analysis (see Figure 1 for an illustration of the pipeline). More concretely, our contributions are as follows.
@@ 1.1 Related work
Our approach is closely related to a line of work that proposes statistical tests for the presence of unobserved confounding. In particular, several works leverage randomized trials to detect unobserved confounding. These tests check for significant differences between average treatment effect estimates obtained from randomized and observational data [viele2014use, yangelastic, morucci2023double, hussain2022falsification]. More sophisticated approaches also test for differences in conditional average treatment effect estimates [hussain2023falsification] and account for right-censored outcomes [demirel2024benchmarking].
Similarly, other works have designed statistical tests using instrumental variables and negative control outcomes instead of randomized trials [lipsitch2010negative, de2014testing, donald2014testing, sofer2016negative]. Additionally, multiple observational studies can be leveraged to test conditional independences and detect unobserved confounding [karlsson2023detecting].
@@
Finally, another line of works proposes calibrating the value of confounding strength using only observational data [hsu2013calibrating, veitch2020sense]. However, the true confounding strength can be arbitrarily different from the calibrated strength, and there is no theoretical result for how these two quantities are related, even with infinite samples.
@@ 2 Setting and notation
where $\mathbb{P}_{\operatorname{det}}$ is given by $Y=Y(T)$, $\mathbb{P}_{\operatorname{inv}}$ is invariant across studies, and $\mathbb{P}_{\operatorname{cnf}}^{\diamond}$ differs for $\diamond\in\{\mathrm{rct},\mathrm{os}\}$. This factorization captures the essence of the potential outcome framework, where $Y(1)$ and $Y(0)$ do not depend on $T$ while being more general. In particular, it allows for shifts in the marginal distribution of the observed and unobserved confounders. We illustrate the corresponding graphical model in Figure 2. While there have been many attempts to unify potential outcomes and graphical models, most prominently Single World Intervention Graphs [richardson2013single], here the graph is purely used for illustrative purposes and does not play a formal role in our analysis.
@@
This property is standard for generalizing the findings of randomized trials to another population [o2014generalizing, colnet2020causal, degtiar2023review], and is a weaker assumption than ignorability of study selection [sugden1984ignorable, hotz2005predicting] or sample ignorability of treatment effects [kern2016assessing].
@@
This assumption is strictly weaker than the positivity of trial participation [stuart2011use, hartman2015sample, andrews2017weighting, nie2021covariate, colnet2022reweighting]. It is also expected to hold in our setting, as it aligns with the design of observational studies by regulatory agencies for drug monitoring  [he2020clinical], and particularly for post-marketing surveillance [franklin2019evaluating, schurman2019framework].
@@ 2.1 Sensitivity analysis
Sensitivity analysis is commonly used to account for unobserved confounding in observational data. In particular, this approach estimates an interval for the treatment effect that depends on an assumed confounding strength $\Gamma$ of $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$. Throughout the paper, we define the confounding strength using the widely accepted marginal sensitivity model [tan2006distributional].
@@
Under this notion of confounding strength, we can define a set of full distributions $\tilde{\mathbb{P}}_{\operatorname{full}}$ that are compatible with the marginal distribution of the observational study $\mathbb{P}^{\operatorname{os}}$ and have a bounded odds ratio.
@@
Given a distribution $\mathbb{P}^{\operatorname{os}}$ over $(X,Y,T)$ and a confounding strength $\Gamma\geq 1$, we define the set $\mathcal{E}(\mathbb{P}^{\operatorname{os}},\Gamma)$ of distributions $\tilde{\mathbb{P}}_{\operatorname{full}}$, as
@@
Above, we do a slight abuse of notation by defining $\mu$ as both a function and a real number, depending on its argument. Several estimators have recently emerged in the literature for the CATE bounds [kallus2019interval, jesson2021quantifying, oprescu2023b] and for the ATE bounds [zhao2019sensitivity, dorn2021doubly, dorn2022sharp]. We will leverage these estimators to construct a statistical test that detects unobserved confounding above a certain strength.
@@ 3.1 Statistical tests for $H_{0}(\Gamma)$
In what follows, we have access to a randomized trial $D_{\operatorname{rct}}=\{(X_{i},Y_{i},T_{i})\}_{i=1}^{n_{\operatorname{rct}}}$ sampled i.i.d. from the distribution $\mathbb{P}^{\operatorname{rct}}$, and an observational study $D_{\mathrm{os}}=\{(X_{i},Y_{i},T_{i})\}_{i=1}^{n_{\operatorname{os}}}$, sampled i.i.d. from the distribution $\mathbb{P}^{\operatorname{os}}$. We first propose estimates for the average treatment effect under two target populations. Then, we leverage these estimates together with the sensitivity bounds to design an asymptotically valid statistical test at significance level $\alpha$. Finally, we show how such a test can be used to establish an asymptotically valid lower bound on the unobserved confounding strength.
@@
$\mu(\mathbb{P}^{\diamond}_{X},\mathbb{P}^{\diamond}_{\operatorname{full}})$ $=\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[Y\left(\frac{T}{\pi}-\frac{1-T}{1-\pi}\right)w(X)\right],\quad\mathrm{where}\quad w(X):=\frac{\mathbb{P}^{\diamond}(X)}{\mathbb{P}^{\operatorname{rct}}(X)}.$
Lemma 3.2 is a well-known result in the transportability literature [cole2010generalizing, colnet2023risk]. It establishes that when the distribution shift between $\mathbb{P}^{\operatorname{rct}}_{X}$ and $\mathbb{P}^{\diamond}_{X}$ can be corrected, we can identify and estimate the ATE under $\mathbb{P}^{\diamond}_{X}$.
@@
  • For $\mathbb{P}^{\diamond}_{X}=\mathbb{P}^{\operatorname{rct}}_{X}$, we estimate the CATE sensitivity bounds from observational data and average them over the target population. Specifically, we use the B-Learner [oprescu2023b] to estimate the sensitivity analysis bounds.
  • For $\mathbb{P}^{\diamond}_{X}=\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}$, we have two options: either estimate the CATE sensitivity analysis bounds and average them, or directly estimate the ATE sensitivity analysis bounds over the target population. In our experiments, we directly estimate the ATE sensitivity analysis bounds using either the DVDS [dorn2021doubly] or the QB estimator [dorn2022sharp].
@@
Fix $\Gamma\in[1,\infty)$ and $\alpha\in(0,1)$, and let $\hat{\phi}_{\diamond}(\Gamma,\alpha)$ be the test defined in Algorithm 1. Under Assumptions 2.12.3, the setting described in Section 2, we have, for $H_{0}(\Gamma)$,
@@
$\hat{\phi}_{\mathrm{rct}}(\Gamma,\alpha)$ is an asymptotically valid level-$\alpha$ test.
(ii) If it holds that $\underset{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}{\lim}n_{\operatorname{rct}}/(n_{\operatorname{rct}}+n_{\operatorname{os}})=\rho\in(0,1)$, $\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}[w(X)^{2}]<\infty$, and $\hat{\mu}_{\Gamma}^{+}$ and $\hat{\mu}_{\Gamma}^{-}$ are consistent estimators of the ATE sensitivity analysis bounds that satisfy
$\sqrt{n_{\operatorname{os}}}\bigl(\hat{\mu}_{\Gamma}^{+}-\mu_{\Gamma}^{+}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}(0,(\sigma_{\Gamma}^{+})^{2}),\qquad\sqrt{n_{\operatorname{os}}}\bigl(\hat{\mu}_{\Gamma}^{-}-\mu_{\Gamma}^{-}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}(0,(\sigma_{\Gamma}^{-})^{2}),$
$\hat{\phi}_{\widetilde{\mathrm{os}}}(\Gamma,\alpha)$ is an asymptotically valid level-$\alpha$ test.
We provide a complete proof in Section A.1.2. Notably, Assumption (ii), together with the fixed limiting proportions of the randomized trial and observational study, is relatively mild and expected to hold for various estimators; for instance, it can be satisfied by the DVDS estimator [dorn2021doubly]. On the other hand, Assumption (i) is stronger and generally only expected to hold when $n_{\operatorname{os}}\gg n_{\operatorname{rct}}$.
In essence, we propose two tests that work under different assumptions: $\hat{\phi}_{\mathrm{rct}}$ relies on a consistent estimate of the CATE sensitivity analysis bounds, while $\hat{\phi}_{\widetilde{\mathrm{os}}}$ relies on directly estimating the ATE sensitivity analysis bounds together with the importance weights $w(X)$. The importance weights can be identified when the observational study and the randomized trial adhere to a nested trial design [olschewski1985comprehensive, olschewski1992analysis, choudhry2017randomized]. See Section A.2 for a discussion on how the importance weights are estimated in this setting.
@@
$\hat{\mu}_{\Gamma}^{+}=\hat{\mathbb{E}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{+}_{\Gamma}(X)],\quad(\hat{\sigma}^{+}_{\Gamma})^{2}=\widehat{\operatorname{Var}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}_{\Gamma}^{+}],\quad\hat{\mu}_{\Gamma}^{-}=\hat{\mathbb{E}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}^{-}_{\Gamma}(X)],\quad(\hat{\sigma}^{-}_{\Gamma})^{2}=\widehat{\operatorname{Var}}_{\mathbb{P}^{\diamond}_{X}}[\hat{\mu}_{\Gamma}^{-}],$
where $\hat{\mathbb{E}}$$\cdot$] denotes the empirical mean, and $\widehat{\operatorname{Var}}$$\cdot$] denotes a consistent estimator of the sampling variance.
@@ 3.2 A lower bound on unobserved confounding strength
Let $\hat{\Gamma}_{\texttt{LB}}$ be as in Equation (2) for a fixed level $\alpha$. Then, under the setting described in Section 2, and the same assumptions as in Proposition 3.1, $\hat{\Gamma}_{\texttt{LB}}$ is an asymptotically valid lower bound,
$\mathbb{P}(\hat{\Gamma}_{\texttt{LB}}\leq\Gamma^{\star})\geq 1-\alpha-o(1).$
@@
$\leq\mathbb{P}(\hat{\phi}_{\mathrm{rct}}(\Gamma^{\star},\alpha)=1)\leq\alpha+o(1),$
@@ 4 Synthetic Experiments
Intuitively, the sensitivity analysis bounds are tight for a specific $\Gamma^{\star}$ when $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}$ leads to a marginal distribution $\mathbb{P}^{\operatorname{os}}$ that maximally biases the estimable ATE. This situation occurs, for instance, when patients experiencing smaller outcomes are assigned to the control group while those with larger outcomes are in the treatment group. In this case, the sensitivity analysis bounds must be sufficiently large to include the true ATE and remain valid. Such a scenario is only possible if $U$ is very informative of $Y(1)$, captured by a high correlation coefficient. Conversely, when the confounder carries little information about the potential outcomes, the sensitivity bounds are unnecessarily conservative, leading to low power of the test and hence looser $\hat{\Gamma}_{\texttt{LB}}$. More formally, Section A.3 discusses two extreme cases. If the unobserved confounder is maximally informative about the potential outcomes, namely $U=(Y(1),Y(0))$, then $\hat{\Gamma}_{\texttt{LB}}$ converges to $\Gamma^{\star}$ in the infinite-sample limit. Conversely, if $U\perp\!\!\!\perp(Y(1),Y(0))$, then the confounder carries no information about the potential outcomes and $\hat{\Gamma}_{\texttt{LB}}$ converges to $1$.